8. A right circular cylinder with an altitude of 40 cm and base diameter of 30cm has a right circular cone at its top with the same base. If the cone is 60cm high, determine its lateral area in cm2.

a) 706.86
b) 2,827.43
c) 2,914.45
d) 3,345.76
1
Expert's answer
2020-12-11T13:13:52-0500

The lateral area of a right circular cone with a slant height "l" and base radius "r" can be found by the formula


"A_L=\pi rl"

By the Pythagorean Theorem from the right triangle


"l^2=r^2+h^2,"

where "h" is an altitude of a cone.

Given "d=30\ cm, h=60\ cm."

Then


"r=\dfrac{d}{2}=\dfrac{30\ cm}{2}=15\ cm"

"l=\sqrt{r^2+h^2}=\sqrt{(15\ cm)^2+(60\ cm)^2}=15\sqrt{17}\ cm"

"A_L=\pi(15\ cm)(15\sqrt{17}\ cm)=(225\sqrt{17})\pi\ cm^2"

"\approx2914.45\ cm^2"

c) 2,914.45